OpenStudy (anonymous):

Xenia is ordering a cake for her friend’s wedding. She knows that there will be 110 guests at the wedding, and she is trying to decide how large a cake to order. She has been told that each guest should get a piece of cake that is 4 inches by 4 inches. She would like to buy a tiered, square three-layer cake and is trying to decide which sizes of cakes to order. Which combination of cakes should she get? The available sizes, in inches, are: 32 × 32 28 × 28 24 × 24 20 × 20 16 × 16 12 × 12 8 × 8

OpenStudy (anonymous):

24*24

OpenStudy (cwrw238):

total area of cake for 110 guests = 110 *4*4 = 1760

OpenStudy (cwrw238):

so e need to pick 3 of the kist to add up to 1760

OpenStudy (anonymous):

since it is three layered , one layer = 1760/3

OpenStudy (anonymous):

ill attach the question .

OpenStudy (anonymous):

the ans. is that option which has a sum of 1760.

OpenStudy (cwrw238):

32^2 = 1024 28^2 = 784 24^2 = 576 20^2 = 400 16^2 =256 12^2 = 144 8^2 = 64 which 3 are = to or nearest and > 1760?

OpenStudy (cwrw238):

how about 2nd + 3rd + 4th?

OpenStudy (anonymous):

alrighty wll i worked it out & i got answer 'C' ! am i right ? !

OpenStudy (anonymous):

*well

OpenStudy (cwrw238):

i can't see your picture but its 28^2 + 24^2 + 20^2

OpenStudy (anonymous):

well then i suppose i got it right because that was answer "C" !

OpenStudy (cwrw238):

yup

OpenStudy (cwrw238):

check it out 784 + 576 + 400 = 1760

OpenStudy (anonymous):

Yes it's right ! (: